# Tidal phenomena in the Galactic Center: The Curious Case of X7

### v1

@@ Abstract
Context. Several enigmatic dusty sources have been detected in the central parsec of the Galactic Center. Among them is X7, located at only 0.02 pc from the central super-massive black hole, Sagittarius A* (Sgr A*). Recent observations have shown that it is becoming elongated due to the tidal forces of Sgr A*. X7 is expected to be fully disrupted during its pericenter passage around 2035 which might impact the accretion rate of Sgr A*. However, its origin and nature are still unknown.
@@
Results. Our results show that ejecta from the star S33/S0-30, launched in 1950, can to a large extent, replicate the observed dynamics and structure of X7, provided that it is initially elongated with a velocity gradient across it, and with an initial maximum speed of 600 km s-1.
Conclusions. Our results show that a grazing collision between the star S33/S0-30 and a field object such as a stellar mass black hole or a Jupiter-mass object is a viable scenario to explain the origin of X7. Nevertheless, such encounters are rare based on the observed stellar dynamics within the central parsec.
@@
The Milky Way Galactic Center (GC) harbors the closest super-massive black hole (SMBH) to Earth: Sagittarius A* (Sgr A*), which has a mass of at a distance of 8.3 kpc (GRAVITY Collaboration et al., 2022). This fact makes it a unique laboratory to study the detailed orbital motion of the stellar and gaseous components in the vicinity of a SMBH (see Genzel et al., 2010; Ciurlo & Morris, 2025, for a review). Over 30 years of near-infrared (NIR) observations have enabled the monitoring of hundreds of stars with high precision (Schödel et al., 2002; Ghez et al., 2003, 2008; Gillessen et al., 2009, 2017; von Fellenberg et al., 2022), and even to test successfully the predictions of General Relativity (e.g. Do et al., 2019a; Amorim et al., 2019; GRAVITY Collaboration et al., 2020). Currently, there are 195 stars with constrained orbits within the central parsec (von Fellenberg et al., 2022). Many of them correspond to B-type stars (Eisenhauer et al., 2005), and tens of O-type and Wolf-Rayet (WR) stars which have significant mass loss in the form of winds (Paumard et al., 2006; Martins et al., 2007; Habibi et al., 2019). Over the last decade, a lot of attention has been put on the dusty sources known as G objects that coexist with the stars in the vicinity of Sgr A* (e.g. Ciurlo et al., 2023). Despite long observation campaigns and theoretical efforts, their nature remains unknown (see Mapelli & Gualandris, 2016; Ciurlo et al., 2020, for a review).
@@
Gillessen et al. (2012) discovered the gaseous and dusty source G2 moving on a highly-eccentric orbit towards Sgr A*. Initially, the source was interpreted as a purely gaseous object with a total mass of merely 3M⊕. Its observed tidal interaction over timescales of the order of years attracted attention due to the potential effect on the quiescent accretion state of Sgr A*. However, there was no clear observed enhancement on the activity of Sgr A* in X-ray (Bouffard et al., 2019) but a bright flare in the near-infrared in 2019 could potentially be attributed to it (Do et al., 2019b; Paugnat et al., 2024). Over the years, many objects of similar characteristics were observed in the central 0.1 pc, increasing the population of G objects up to ten (e.g. Pfuhl et al., 2015; Witzel et al., 2017; Ciurlo et al., 2020; Peißker et al., 2020, 2024b). Many of them show extended emission of the Br recombination line at 2.1661 m and point-source emission in the L′ band at 3.776 m. None of them show obvious hints for the presence of a stellar object within the extended source. As a result, theoretical efforts have focused on constraining their true nature, especially of G2. But no consensus has been reached on its nature and origin yet. The main debate is focused on whether G2 is a purely gaseous cloud or harbors a compact (stellar) object. In the former, it has been suggested that G2 could be a gas clump formed in stellar wind collisions (Burkert et al., 2012; Calderón et al., 2016), the result of a slow wind from a Luminous Blue Variable (LBV) star (Burkert et al., 2012), or a nova outburst from the partial tidal disruption of a giant star (Meyer & Meyer-Hofmeister, 2012), among others. In the latter, G2 has been hypothesized as an evaporating circumstellar disk (Miralda-Escudé, 2012; Murray-Clay & Loeb, 2012), a proto-planet (Mapelli & Ripamonti, 2015), a mass-losing low-mass star (Ballone et al., 2013; De Colle et al., 2014; Valencia-S. et al., 2015; Ballone et al., 2016, 2018), the product of a stellar merger (Witzel et al., 2014; Prodan et al., 2015; Ciurlo et al., 2020), or a young stellar object (YSO; Peißker et al., 2024b). Moreover, Peißker et al. (2024a) also argue for binaries within the G object population.
@@
In this region at a separation of 0.5″(0.02 pc) from Sgr A*, the source X7 has been observed to be interacting tidally with the central black hole. Although X7 has certain similarities with the G objects, e.g. it is an extended source made out of gas and dust, it is much larger, reaching 3300 au in 2021 whereas G-objects are of the order of 100 au. It was reported for the first time by Clénet et al. (2004) and since then it has been routinely monitored by several groups (Mužić et al., 2007; Peißker et al., 2020, 2021; Ciurlo et al., 2023). Initially, it was thought to be a bow shock but recent observations suggest a morphological deviation from that picture (Ciurlo et al., 2023). Recently, Peißker et al. (2021) proposed that X7 is a circumstellar envelope of the star S50/S0-73 as they were close in earlier observations. However, Ciurlo et al. (2023) argued against that hypothesis by revealing that there is a significant three-dimensional spatial and dynamical separation between them. Additionally, they calculated that the pericenter passage of X7 is estimated to take place around 2035. If X7 is a purely gaseous and dusty source, we will be able to witness the tidal disruption of this object by Sgr A* and consequently enhanced accretion activity of Sgr A*. An alternative scenario has been proposed recently by Peißker et al. (2024b), which suggests both X7 and a similar dusty feature called X3 correspond to YSOs. Nevertheless, the nature of X7 is still unknown and observational and theoretical efforts are ongoing to unravel its mystery.
@@
In this work, we present a study to constrain the origin of X7. We work under the assumption that the source is purely gaseous and dusty, and therefore its age is less than its orbital period of 200 years. We present a set of test-particle simulations and analytical estimates to study how likely it is that X7 was related to any of the stars in the central parsec of the GC. Our results show that the current observations of X7 are consistent with the source being the ejecta resulting from a grazing collision between the S-star S33/S0-30 (an early B-type star) and a field object such as a stellar-mass black hole or a Jupiter-mass object. It has been suggested that X7 could be connected to the G object G3 (Ciurlo et al., 2023), because of the remarkable alignment of their orbits. However, we find that they are most likely not connected since, provided the age of X7 to be less than 200 yrs, they show significantly different orbital phases even considering the large associated uncertainties. This article is organized as follows: in Section 2, we present the properties of X7 as well as the three hypotheses for its formation studied in this work. Section 3 contains the details of the calculations, methods and the numerical setup for testing the hypotheses. In Section 4, we present the results from our analysis and simulations. Section 5 discusses the implication of our results for understanding the origin of X7. Finally, we summarize and conclude this study in Section 6. Throughout this paper, we used and , where and are the mass of Sgr A* and the distance to it from Earth (GRAVITY Collaboration et al., 2022).
@@ 2 Hypotheses for the origin of X7
X7 has been monitored continuously since 2002 with telescopes at the W. M. Keck Observatory (Ciurlo et al., 2023) as part of the Galactic Center Orbits Initiative (GCOI; PI: Andrea Ghez). The data traced the dust thermal emission – through images in L′-band (3.776 m) obtained with the imager NIRC2 - and the gas emission through the Br- hydrogen recombination line observed with the integral field spectrograph OSIRIS (Larkin et al., 2006). These observations have shown that the size of X7 has nearly doubled from au in 2003 to au in 2021. Additionally, it was possible to fit a Keplerian orbit to its forward tip, revealing a semi-major axis of au and an eccentricity of (Ciurlo et al., 2023). Under the assumption that it is a purely gas and dust feature, a total mass of 50 M⊕ was derived using the observed Br- flux. Then, it is straightforward to estimate that the ratio between the derived and Roche densities (the critical density to remain intact against tidal forces of the SMBH) of X7 is very small (10-5). This shows that its self-gravity is negligible provided that no compact source is associated to it. Based on this, its age is constrained to be less than its orbital period (200 yr), as otherwise the tidal forces would have already destroyed it.
@@
First, X7 could be the result of a mass outflow from a massive star going through a high mass-loss episode such as a Luminous Blue Variable (LBV) phase. This scenario is plausible since the central parsec hosts hundreds of O- and B-type stars (e.g. Paumard et al., 2006; Martins et al., 2007; Habibi et al., 2019).
The second hypothesis conceives X7 as ejecta from a binary merger process. This case is supported by the fact that the sources G3 and X7 could be considered dynamically linked as they move on very similar orbits and share common observational properties (see Ciurlo et al., 2020, 2023). In this scenario, G3 would be the merger product and X7 the mass ejecta from the Eccentric Kozai-Lidov (EKL; Kozai, 1962; Lidov, 1962; Naoz, 2016) induced merger process.
@@
Schematic representations of the hypotheses are shown in Figure 1.
@@ 3 Analysis
Regardless of the hypothesis considered, all of them have in common that X7 would be related to a stellar source in the region. Hence, as a first step we proceed to investigate which stars could be linked to X7 based on their dynamics. It is to be noted that, when we mention the orbit of X7, we refer to the orbit of its tip (Ciurlo et al., 2023). Once we find the best star candidate to be related to X7, we simulate the dynamical evolution of a cloud of test particles launched from the position and time of the closest encounter. In this section, we describe in detail these calculations.
@@ 3.1 Correlation with stars in the Galactic Center
First, we analyzed whether X7 is related to one of the stars in the vicinity of Sgr A*. Thus, we searched for stars that were close simultaneously in space and time from X7 by tracing their orbits back in time into the past. To do so, we assumed that the orbits have not changed significantly in the last 200 years. This is a reasonable assumption since the dominant gravitational field is due to Sgr A*. Currently, there are 36 stars with fully constrained orbital parameters and 159 stars with constrained orbits but incomplete orbital solutions (-coordinate missing) (von Fellenberg et al., 2022). Following the MINECC method in Cuadra et al. (2008), we looked for the minimum eccentricity possible for each of the 159 stars and considered the corresponding value of in order to model their orbits. By doing so, we have a total of 195 stars with orbits for our analysis.
@@
As a next step, we inspected the angular momentum direction of the candidate stars and X7. Figure 3 is a Hammer projection that shows the orbital angular momentum direction for all stars in our sample as well as the sources X7 and G3 (Ciurlo et al., 2020, 2023). The stars with completely determined orbits are shown as orange star symbols while the rest of the 159 are displayed as sky-blue star symbols. S14/S0-16, S33/S0-30, X7, and G3 are highlighted as teal, magenta, blue, and green star symbols, respectively. Notice that S14/S0-16 has a very different angular momentum orientation compared to both X7 and G3, whereas S33/S0-30 is much closer to X7 and G3 in the Hammer projection. In principle, this points to S33/S0-30 being a more promising candidate for being related to X7. Despite some stars having similar angular momentum orientation to X7, it is important to bear in mind that they have a large minimum three-dimensional separation to X7 ( au), so they are unlikely to be related. Gillessen et al. (2009, 2017) have characterized the star S33/S0-30 as an early-type star that is part of the S-star cluster. Unfortunately, we are not aware of additional information on this star.
@@ 3.2 Test-particle simulations
We model the evolution of a cloud represented with test particles launched from the position and time of the closest encounter between S33/S0-30 and X7. First, we describe the numerical setup and technique. Then, we calculate the initial conditions for the simulations such as initial velocity, direction, radius, and shape of the cloud.
@@ 3.2.1 Numerical Setup
where is the radial distance of the particle to the location of Sgr A*, is the mass of the SMBH, is the gravitational constant, and is the prefactor which is set to for accuracy. We have applied an adaptive time-stepping scheme where the minimum timestep out of all the particles at each iteration is considered. The initial position of the cloud was set to the position of S33/S0-30 in 1947. Then, it is only necessary to specify the size of the cloud and its initial velocity.
@@ 3.2.2 Initial velocity
To find the initial velocity of the cloud, we calculated the velocity required for ejecta from the position of S33/S0-30 in 1947 to achieve the same observed specific angular momentum of the tip of X7. To do so, we solved the following equation to find ,
This corresponds to a system of linear equations obtained that is not independent, so a unique solution is not possible. However, for each value of there are unique values of and . The components and can be obtained from the below equations,
Then, we sampled a wide range of values for from 10000 km s-1 and 10000 km s-1 with a step size of 1 . In this way, the velocity magnitude will be of the same order as the typical velocities of the stars in the GC. To select the optimal initial velocity and its direction, we selected the velocities that minimize the three-dimensional separation and relative 3D velocity between the tip of the simulated cloud and the observed tip of X7 in 2021, which are represented as and , respectively. Figure 4 shows and as a function of the magnitude of the initial velocity in 1947. It is important to remark that only a zoomed version of the region around the minima is shown. Here it is possible to observe that an initial velocity of minimizes both the and .
For a range of , between , the optimal range of components of initial velocity, having an initial velocity magnitude of , is between
By optimal, we mean that the separation of the tip of the simulated cloud and the observed tip of X7 () is less than 175 au (21 mas), which falls within the margin of error in the 2021 observations from Ciurlo et al. (2023).

### v2

@@ Abstract
Context. Several enigmatic dusty sources have been detected in the central parsec of the Galactic Center. Among them is X7, located at only 0.02 pc from the central supermassive black hole, Sagittarius A* (Sgr A*). Recent observations have shown that X7 is becoming elongated due to the tidal forces of Sgr A*. X7 is expected to be fully disrupted during its pericenter passage around 2035, which might impact the accretion rate of Sgr A*. However, its origin and nature are still unknown.
@@
Results. Our results show that ejecta from the star S33/S0-30, launched in 1950, can to a large extent replicate the observed dynamics and structure of X7, provided that it is initially elongated with a velocity gradient across it, and with an initial maximum speed of 600 km s-1.
Conclusions. Our results show that a grazing collision between the star S33/S0-30 and a field object such as a stellar-mass black hole or a Jupiter-mass object is a viable scenario to explain the origin of X7. Despite the uncertainties in the rate of these encounters, recent estimations show that it is plausible for such a scenario to have occurred recently.
@@
The Milky Way Galactic Center (GC) harbors the closest supermassive black hole (SMBH) to Earth: Sagittarius A* (Sgr A*), which has a mass of at a distance of 8.3 kpc (GRAVITY Collaboration et al., 2022). This makes it a unique laboratory to study the detailed orbital motion of the stellar and gaseous components in the vicinity of a SMBH (see Genzel et al., 2010; Ciurlo & Morris, 2025, for a review). Over 30 years of near-infrared (NIR) observations have enabled the monitoring of hundreds of stars with high precision (Schödel et al., 2002; Ghez et al., 2003, 2008; Gillessen et al., 2009, 2017; von Fellenberg et al., 2022), and even the successful testing of the predictions of general relativity (e.g., Do et al., 2019a; Amorim et al., 2019; GRAVITY Collaboration et al., 2020). Currently, there are 195 stars with constrained orbits within the central parsec (von Fellenberg et al., 2022). Many of them are B-type stars (Eisenhauer et al., 2005); there are also tens of O-type and Wolf-Rayet (WR) stars that have significant mass loss in the form of winds (Paumard et al., 2006; Martins et al., 2007; Habibi et al., 2019). Over the last decade, a lot of attention has been paid to the dusty sources known as G objects that coexist with the stars in the vicinity of Sgr A* (e.g., Ciurlo et al., 2023). Despite long observation campaigns and theoretical efforts, their nature remains unknown (see Mapelli & Gualandris, 2016; Ciurlo et al., 2020, for a review).
@@
Gillessen et al. (2012) discovered the gaseous and dusty source G2, moving on a highly eccentric orbit toward Sgr A*. Initially, the source was interpreted as a purely gaseous object with a total mass of merely 3M⊕. Its observed tidal interaction over timescales on the order of years attracted attention due to its potential effect on the quiescent accretion state of Sgr A*. However, there was no clear observed enhancement in the activity of Sgr A* in X-rays (Bouffard et al., 2019), but a bright flare in the NIR in 2019 could potentially be attributed to it (Do et al., 2019b; Paugnat et al., 2024). Over the years, many objects with similar characteristics have been observed in the central 0.1 pc, increasing the known population of G objects to ten (e.g., Pfuhl et al., 2015; Witzel et al., 2017; Ciurlo et al., 2020; Peißker et al., 2020, 2024b). Many show extended emission of the Brackett gamma (Br) recombination line at 2.1661 m and point-source emission in the L′ band at 3.776 m. None show obvious hints of a stellar object within the extended source. As a result, theoretical efforts have focused on constraining their true nature, especially that of G2, but no consensus has been reached on its nature and origin. The main debate is focused on whether G2 is a purely gaseous cloud or harbors a compact (stellar) object. Regarding the purely gaseous cloud hypothesis, it has been suggested that G2 could be a gas clump formed in stellar wind collisions (Burkert et al., 2012; Calderón et al., 2016), the result of a slow wind from a luminous blue variable (LBV) star (Burkert et al., 2012), or a nova outburst from the partial tidal disruption of a giant star (Meyer & Meyer-Hofmeister, 2012), among others. Within the compact object argument, G2 has been hypothesized to be an evaporating circumstellar disk (Miralda-Escudé, 2012; Murray-Clay & Loeb, 2012), a protoplanet (Mapelli & Ripamonti, 2015), a mass-losing low-mass star (Ballone et al., 2013; De Colle et al., 2014; Valencia-S. et al., 2015; Ballone et al., 2016, 2018), the product of a stellar merger (Witzel et al., 2014; Prodan et al., 2015; Ciurlo et al., 2020), or a young stellar object (Peißker et al., 2024b). Moreover, Peißker et al. (2024a) also argue that there are binaries within the G object population.
@@
In this region, at a separation of 0.5″(0.02 pc) from Sgr A*, the source X7 has been observed to be interacting tidally with the central black hole. Although X7 has certain similarities with the G objects, for example it is an extended source composed of gas and dust, it is much larger, reaching 3300 au in 2021, whereas G objects are on the order of 100 au. It was reported for the first time by Clénet et al. (2004) and since then has been routinely monitored by several groups (Mužić et al., 2007; Peißker et al., 2020, 2021; Ciurlo et al., 2023). Initially, it was thought to be a bow shock, but recent observations suggest a morphological deviation from that picture (Ciurlo et al., 2023). Peißker et al. (2021) propose that X7 is a circumstellar envelope of the star S50/S0-73 since they were close in earlier observations. However, Ciurlo et al. (2023) argue against that hypothesis; they find that there is a significant three-dimensional spatial and dynamical separation between them. Additionally, they estimate that the pericenter passage of X7 will take place around 2035. If X7 is a purely gaseous and dusty source, we will be able to witness the tidal disruption of this object by Sgr A* and, consequently, the enhanced accretion activity of Sgr A*. An alternative scenario has recently been proposed by Peißker et al. (2024b), who suggest that both X7 and a similar dusty feature called X3 are young stellar objects. Nevertheless, the nature of X7 is still unknown, and observational and theoretical efforts are ongoing to unravel its mystery.
@@
We have carried out a study to constrain the origin of X7. We worked under the assumption that the source is purely gaseous and dusty and, therefore, that its age is less than its orbital period of 200 years. We present a set of test-particle simulations and analytical estimates to study how likely it is that X7 was related to any of the stars in the central parsec of the GC. Our results show that the current observations of X7 are consistent with the source being the ejecta resulting from a grazing collision between the S star S33/S0-30 (an early B-type star) and a field object such as a stellar-mass black hole or a Jupiter-mass object. It has been suggested that X7 could be connected to the G object G3 (Ciurlo et al., 2023) because of the remarkable alignment of their orbits. However, we find that they are most likely not connected since, provided the age of X7 is less than 200 yrs, they show significantly different orbital phases, even when considering the large associated uncertainties. ?
This article is organized as follows: In Sect. 2 we present the properties of X7 as well as the three hypotheses for its formation studied in this work. Section 3 contains the details of the calculations, the methods, and the numerical setup for testing the hypotheses. In Sect. 4 we present the results from our analysis and simulations. Section 5 discusses the implication of our results for understanding the origin of X7. Finally, we summarize and conclude this study in Sect. 6. Throughout this paper, we use and , where and are the mass of Sgr A* and the distance to it from Earth (GRAVITY Collaboration et al., 2022).
@@ 2 Hypotheses for the origin of X7
X7 has been monitored continuously since 2002 with telescopes at the W.M. Keck Observatory (Ciurlo et al., 2023) as part of the Galactic Center Orbits Initiative (GCOI; PI: Andrea Ghez). The data traced the dust thermal emission – through images in the L′ band (3.776 m) obtained with the imager NIRC2 – and the gas emission through the Br hydrogen recombination line observed with the integral field spectrograph OSIRIS (Larkin et al., 2006). These observations have shown that the size of X7 has nearly doubled from au in 2003 to au in 2021. Additionally, it was possible to fit a Keplerian orbit to its forward tip, revealing a semimajor axis of au and an eccentricity of (Ciurlo et al., 2023). Under the assumption that it is a purely gas and dust feature, a total mass of 50 M⊕ was derived using the observed Br flux. Then, it is straightforward to estimate that the ratio between the derived and Roche densities (the critical density to remain intact against tidal forces of the SMBH) of X7 is very small (10-5). This shows that its self-gravity is negligible provided that no compact source is associated with it. Based on this, its age is constrained to be less than its orbital period (200 yr), as otherwise the tidal forces would have already destroyed it.
@@
The second hypothesis conceives X7 as ejecta from a binary merger process. This case is supported by the fact that the sources G3 and X7 could be considered dynamically linked as they move on very similar orbits and share common observational properties (see Ciurlo et al., 2020, 2023). In this scenario, G3 would be the merger product and X7 the mass ejecta from the eccentric Kozai-Lidov (EKL)-induced merger process (Kozai, 1962; Lidov, 1962; Naoz, 2016).
@@
Schematic representations of the hypotheses are shown in Fig. 1.
@@ 3 Analysis
Regardless of the hypothesis considered, all of them have in common that X7 would be related to a stellar source in the region. Hence, as a first step we investigated which stars could be linked to X7 based on their dynamics. It is to be noted that when we mention the orbit of X7, we are referring to the orbit of its tip (Ciurlo et al., 2023). Once found find the best star candidate to be related to X7, we simulated the dynamical evolution of a cloud of test particles launched from the position and time of the closest encounter. In this section we describe in detail these calculations.
@@ 3.1 Correlation with stars in the Galactic Center
First, we analyzed whether X7 is related to one of the stars in the vicinity of Sgr A*. Thus, we searched for stars that were close simultaneously in space and time from X7 by tracing their orbits back in time into the past. To do so, we assumed that the orbits have not changed significantly in the last 200 years. This is a reasonable assumption since the dominant gravitational field is due to Sgr A*. Currently, there are 36 stars with fully constrained orbital parameters and 159 stars with constrained orbits but incomplete orbital solutions (-coordinate missing; von Fellenberg et al. 2022). Following the MINECC method in Cuadra et al. (2008), we looked for the minimum eccentricity possible for each of the 159 stars and considered the corresponding value of in order to model their orbits. By doing so, we have a total of 195 stars with orbits for our analysis.
@@
As a next step, we inspected the angular momentum direction of the candidate stars and X7. Figure 3 is a Hammer projection that shows the orbital angular momentum direction for all stars in our sample as well as the sources X7 and G3 (Ciurlo et al., 2020, 2023). Notice that S14/S0-16 has a very different angular momentum orientation compared to both X7 and G3, whereas S33/S0-30 is much closer to X7 and G3 in the Hammer projection. In principle, this indicates that S33/S0-30 is a more promising candidate for being related to X7. Despite some stars having a similar angular momentum orientation to that of X7, it is important to bear in mind that they have a large minimum three-dimensional separation to X7 ( au), so they are unlikely to be related. Gillessen et al. (2009, 2017) have characterized the star S33/S0-30 as an early-type star that is part of the S-star cluster. Unfortunately, we are not aware of additional information on this star.
@@ 3.2 Test-particle simulations
We modeled the evolution of a cloud represented with test particles launched from the position and time of the closest encounter between S33/S0-30 and X7. Here we describe the numerical setup and technique. Then, we show how we calculated the initial conditions for the simulations, such as the initial velocity, direction, radius, and shape of the cloud.
@@ 3.2.1 Numerical setup
where is the radial distance of the particle to the location of Sgr A*, is the mass of the SMBH, is the gravitational constant, and is the prefactor, which is set to for accuracy. We applied an adaptive time-stepping scheme in which the minimum timestep, , of all the particles at each iteration is considered. The initial position of the cloud was set to the position of S33/S0-30 in 1947. Then, it is only necessary to specify the size of the cloud and its initial velocity.
@@ 3.2.2 Initial velocity
To find the initial velocity of the cloud, we calculated the velocity required for ejecta from the position () of S33/S0-30 in 1947 to achieve the same observed specific angular momentum () of the tip of X7. To do so, we solved the following equation to find ,
This corresponds to a system of linear equations obtained that is not independent, so a unique solution is not possible. However, for each value of there are unique values of and . The components and can be obtained from the following equations:
Then, we sampled a wide range of values for from 10000 km s-1 and 10000 km s-1 with a step size of 1 . In this way, the velocity magnitude will be of the same order as the typical velocities of the stars in the GC. To select the optimal initial velocity and its direction, we selected the velocities that minimize the three-dimensional separation and relative three-dimensional velocity between the tip of the simulated cloud and the observed tip of X7 in 2021, which are represented as and , respectively. Figure 4 shows and as a function of the magnitude of the initial velocity () in 1947. It is important to remark that only a zoomed version of the region around the minima is shown. Here it is possible to observe that an initial velocity of minimizes both the and .
For a range of , , the optimal range of components of initial velocity, having an initial velocity magnitude of , is between
By ”optimal,” we mean that the separation of the tip of the simulated cloud and the observed tip of X7 () is less than 175 au (21 mas), which falls within the margin of error in the 2021 observations from Ciurlo et al. (2023).